​(Complex annuity​)Upon graduating from college

35 years​ ago, Dr. Nick Riviera was already thinking of retirement. Since​ then, he has made deposits into his retirement fund on a semiannual basis in the amount of ​$700.

Nick has just completed his final payment and is at last ready to retire. His retirement fund has earned

11.8​%

interest compounded semiannually.semiannually.

a. How much has Nick accumulated in his retirement​ account?

b. In addition to all​ this,15 years​ ago, Nick received an inheritance check for

​$25,000 from his beloved uncle. He decided to deposit the entire amount into his retirement fund. What is his current balance in the​ fund?

a. Since graduating from college35

years​ ago, Dr. Nick Riviera has made deposits into his retirement fund on a

semiannual basis in the amount of $700.

If his retirement fund has earned

11.8​%

interest compounded

semiannuallysemiannually​,

how much has Nick accumulated in his retirement​ account?
Answer #1

a) Amount accumulated can be calculated using FV function

N = 35 x 2 = 70, I/Y = 11.8%/2 = 5.9%, PMT = 700, PV = 0

=> Compute FV = $644,225.32

b) Value of inheritance at retirement, FV = PV x (1 + r)^n = 25,000 x (1 + 5.9%)^(15 x 2) = $139,578.59

Total Value = $783,803.91